OpenStudy (anonymous):

What is the slope of the line passing through (1, 2) and (3, 8)? a.) slope = 1/7 slope = 1/3 b.) slope = 1/3 c.) slope = 3 d.) slope = 7

OpenStudy (trojanpoem):

y_2 - y_1 / x_2 - x_1 = 8 -2 / 3- 1 = 6/ 2 = 3